OpenStudy (korosh23):

Pre-calculus 12 Question! Logarithm function convert to exponential function! Convert this to exponential form: y= 2 log3 (2x)

OpenStudy (anonymous):

you gottta bring in the two first

OpenStudy (anonymous):

i.e. start with \[y=\log_3(4x^2)\]

OpenStudy (korosh23):

this is what I think: 3^y/2 = 2x

OpenStudy (anonymous):

then \[y=\log_b(x)\iff b^y=x\]

OpenStudy (korosh23):

oh I did not know you can times the 2 out side with inside 2x?

OpenStudy (anonymous):

you could write it your way, yes

OpenStudy (anonymous):

\[y=2\log_3(2x)\\ \frac{y}{2}=\log_3(2x)\\ 3^{\frac{y}{2}}=2x\]

OpenStudy (korosh23):

I see

OpenStudy (anonymous):

which is the same as \[3^y=(2x)^2\\ 3^y=4x^2\]

OpenStudy (korosh23):

yes

OpenStudy (anonymous):

i used this first \[n\log_b(x)=\log_b(x^n)\]

OpenStudy (anonymous):

get the same answer of course

OpenStudy (korosh23):

just a quick check: I looked up on desmos, y= log3 (4x) does not equal y = 2 log 3 (2x)

OpenStudy (korosh23):

@satellite73

OpenStudy (anonymous):

no it doesn't

OpenStudy (anonymous):

\[2\log(2x)=\log((2x)^2)=\log(4x^2)\]

OpenStudy (korosh23):

including base 3?

OpenStudy (anonymous):

including base your favourite number base is irrelevant, it is a property of any log

OpenStudy (korosh23):

ok got it. Thank you :)

OpenStudy (anonymous):

yw